how many milliliters of a 15% alcohol solution must be added to 80 mL of an 8% alcohol solution to make a 12% alcohol solution?
Answer by stanbon(75887)

how many milliliters of a 15% alcohol solution must be added to 80 mL of an 8% alcohol solution to make a 12% alcohol solution?
------
Equation:
alc + alc = alc
0.15x + 0.08*80 = 0.12(x + 80)
--------------------------------
15x + 8*80 = 12x + 12*80
3x = 4*80
x = 106 2/3 ml (amt. of 15% solution needed)
=========================
Cheers,
Stan H.
=========================